# CroCoDai: A Stablecoin for Cross-Chain Commerce

### v1

@@
Abstract
Decentralized Finance (DeFi), in which digital assets are exchanged without trusted intermediaries, has grown rapidly in value in recent years. The global DeFi ecosystem is fragmented into multiple blockchains, fueling the demand for cross-chain commerce. Existing approaches for cross-chain transactions, e.g., bridges and cross-chain deals, achieve atomicity by locking assets in escrow. However, locking up assets increases the financial risks for the participants, especially due to price fluctuations and the long latency of cross-chain transactions. Stablecoins, which are pegged to a non-volatile asset such as the US dollar, help mitigate the risk associated with price fluctuations. However, existing stablecoin designs are tied to individual blockchain platforms, and trusted parties or complex protocols are needed to exchange stablecoin tokens between blockchains.
Our goal is to design a practical stablecoin for cross-chain commerce. Realizing this goal requires addressing two challenges. The first challenge is to support a large and growing number of blockchains efficiently. The second challenge is to be resilient to price fluctuations and blockchain platform failures. We present CroCoDai to address these challenges. The key insight is to extend the existing Dai stablecoin system to a stablecoin that is backed with collateral tokens from multiple blockchains. CroCoDai leverages the system’s governance layer (which could be another blockchain) to exchange tokens between different blockchains – this enables secure cross-chain commerce and addresses the first challenge. Furthermore, the relatively low correlation between the price of assets on different blockchains makes CroCoDai more resilient to financial shocks than a single-chain stablecoin. We demonstrate this advantage through an empirical analysis that uses real token price data from a period that includes the collapse of the FTX exchange, which caused high token price volatility. We present three prototype implementations of our stablecoin system, and show that it incurs small execution overhead.
keywords:
Stablecoins, blockchains, interoperability, financial risk analysis
category:
@@ Abstract
1 Introduction
Decentralized Finance (DeFi) promises to revolutionize finance by removing the need for trusted middlemen to process transactions. In DeFi, transactions commonly involve a change of ownership of digital tokens that exist on a distributed ledger such as a blockchain – e.g., one user sending bitcoins to another [45]. Many blockchain platforms (e.g., Ethereum [8]) also support smart contracts, which enable automated versions of financial concepts such as auctions, loans, swaps, decentralized exchanges, and ownership shares. The total value of the tokens held in DeFi smart contracts peaked at over \$150 billion USD in late 2021 [53, 15], with at least \$100 billion held in decentralized exchanges alone [55], and has since fallen to around \$45-50 billion as of June 2023 [15].
The DeFi ecosystem is currently fragmented over a multitude of blockchains which each offer their own choices in terms of efficiency, security, and functionality. For example, Cardano [28, 31], Solana [57], Avalanche [49], and Algorand [24] each had a valuation of more than a billion USD in early 2023. These blockchains are not designed to interoperate: i.e., there is no native mechanisms for transferring tokens between chains. However, the recent emergence of blockchain bridges [19] and dedicated interoperability platforms such as Cosmos [7] and Polkadot [33] demonstrates an increasing demand for cross-chain interoperability. A major challenge in cross-chain trades is atomicity: the transactions in a trade must be committed on all blockchains, or not at all. For example, in a cross-chain auction, the first party to transfer tokens (e.g., the auctioneer) cannot be sure that the other party will uphold their part of the trade in the absence of a cross-chain protocol. This is of particular concern in DeFi, as individual blockchain transactions are irreversible by design.
Existing solutions that provide cross-chain atomicity, including relays and hashed time locks [27, 2, 50], essentially lock the assets involved in a trade and only release (or return) them once the transactions in all of the blockchains have been committed (or aborted). While ensuring atomicity, this approach increases the financial risks for the users in the following ways: 1. the tokens held in escrow may lose value due to price fluctuations, 2. one party may abort the transaction, leading to an opportunity cost for the other parties, and 3. the DeFi platforms that facilitate the trade may be faulty or malicious . These risks are interrelated: if there is a significant change in the value of the tokens in escrow (1.), then this may cause a party to abandon a trade (2.), creating an opportunity cost to the other users. The risks identified above are a limitation of all cross-chain protocols that hold tokens in escrow – the longer the tokens are held, the longer they are exposed to DeFi’s notorious price volatility [1] and vulnerability to attacks [60]. Although the abandonment of trades can be disincentivized through the use of premiums [56], price fluctuations can be mitigated in principle if the tokens held in escrow are stablecoins [32, 9, 44] whose value is pegged to a less volatile real-world asset (e.g., the US dollar). However, a flawed stablecoin design can itself put users’ tokens at risk, as witnessed by the collapse of the Terra/Luna stablecoin in May 2022 [6], in which stablecoins worth 60 billion USD lost nearly all of their value.
Our goal is to design a practical stablecoin as a central point of exchange for cross-chain commerce. Realizing this goal requires that two main challenges are addressed. The first challenge is to support a large and growing number of blockchains efficiently, including new and relatively low-use platforms. For established platforms, Automated Market Makers (AMMs, [17]) such as Uniswap can be used to exchange tokens for stablecoins within minutes, and the stablecoins can be transferred between blockchains using a wallet on a cryptocurrency exchange such as Coinbase or Binance. However, the latter step may take several hours before the withdrawal on the receiving chain is processed (e.g., 2-3 hours on Crypto.com [12]), is subject to the exchange’s transaction fees, and the exchange is a single point of failure. Furthermore, AMMs depend on depositors to deposit collateral for each supported pair of cryptocurrencies, and this may not be available for smaller platforms – similarly, centralized exchanges may not always support transfers on smaller platforms. Just as an AMM requires depositors willing to create liquidity pairs, Xclaim [58] requires the availability of a custodian who is willing to put up collateral on the receiving chain, which can be a challenge for small blockchains. Finally, secure bridges such as zkBridge [54] either require a direct bridge to transfer stablecoins between two small blockchains, which may again be unavailable, or require that they are sent via a larger blockchain in a two-step process, which incurs additional time and transaction fee costs. The second challenge in designing a cross-chain stablecoin is to ensure that it is resilient to black swan events (e.g., 51% attacks) that affect individual blockchain platforms, and to price fluctuations that affect multiple blockchains simultaneously. This is particularly relevant if the stablecoin’s peg is maintained through tokens held as collateral on different blockchains. If the total value of the collateral drops below the intended value of the issued stablecoins, then the stablecoin may lose its peg, forcing holders to sell their tokens at a loss – we will refer to such an event as system failure.
In this work, we present CroCoDai, a stablecoin for cross-chain commerce that addresses the above challenges. The key insight is to back CroCoDai with collateral tokens from a multitude of blockchains, using the same principles as the existing Dai stablecoin system (DSS) [38] but without being restricted to the Ethereum ecosystem. To that end, CroCoDai manages communication between the different blockchains through a governance layer which can be implemented through a quorum of trusted nodes, or on top of an existing blockchain. By using the governance layer to create/destroy the same amount of stablecoins on different chains, CroCoDai supports cross-chain transfers by design. Crucially, CroCoDai allows tokens on smaller blockchains to be backed by collateral on larger chains, thus providing stablecoin support to emerging platforms and addressing the first challenge. Unlike Polkadot or Cosmos, CroCoDai achieves its goal without requiring intrusive changes to existing blockchains. To demonstrate how well CroCoDai addresses the second challenge, we present a technical analysis that shows that attacks on a limited number of blockchain platforms can be tolerated as long as the system is sufficiently overcollateralized. We also show that attacks against a limited number of price oracles (which determine the value of the collateral) can be tolerated – this is not a challenge specific to multi-chain settings, so this security property holds for the single-chain DSS as well. Next, we provide evidence for the greater resilience of a multi-chain stablecoin to price fluctuations through an empirical analysis based on the collected data from several periods, including the collapse of the FTX exchange which led to severe depreciation of the Solana cryptocurrency. Specifically, we compare (1) a portfolio that approximates the current state of the DSS’ collateral to a similar portfolio with tokens from multiple chains, and (2) a portfolio that is optimized for low failure risk using the tokens supported by the DSS to one with tokens from multiple chains. Finally, we provide a reference implementation for CroCoDai that uses a dedicated relay chain, and use experiments to show that the computational overhead of CroCoDai is small. The code/data used in this work have been temporarily redacted to preserve anonymity, but they will be open-sourced at a later date to facilitate reproducibility and reusability.
Contributions
Our work makes the following contributions:
  • We present CroCoDai, a cross-chain stablecoin – we present two design variants: one uses a quorum of nodes as a trust anchor; the other uses an existing blockchain (Section 4).
  • We demonstrate CroCoDai’s security through a technical analysis that demonstrates that CroCoDai can tolerate a limited number of compromised blockchains or price oracles (Section 5.1), and an empirical analysis in which we compare single-chain portfolios to portfolios that use collateral from multiple chains, and find that the latter has lower failure probabilities. Specifically, the failure probability is roughly a third lower for optimized pure-crypto portfolios (Sections 5.2 and 5.3).
  • We present three implementations of CroCoDai subvariants that use the quorum-based approach and compare their gas costs and processing times, to demonstrate CroCoDai’s practicality and to make its generality/performance trade-offs explicit (Section 6).
@@ 1 Introduction
The remainder of the paper is structured as follows. Section 2 provides background on DeFi and related work. Section 3 presents the system model. Section 4 presents the design of CroCoDai. Section 5 contains the analysis of CroCoDai’s resilience against black swan events and price fluctuations. Section 6 details the three subvariant implementations and corresponding experimental results. Section 7 discusses future work and concludes the paper.
2 Background
2.1 General Background
Blockchains & Tokens. DeFi uses cryptographically secured, decentralized data structures to replace trusted intermediaries. Blockchains are by far the most prominent example of such data structures. A blockchain is an ordered list of transactions grouped into blocks. The principal function of transactions in DeFi is to transfer tokens, which represent ownership of assets. There are two common types of tokens: fungible tokens – also commonly called cryptocurrencies – and non-fungible tokens (NFTs) [13]. Fungible tokens of the same type are identical, whereas each NFT represents a unique asset.
In the first cryptocurrency platform, Bitcoin [45], each transaction transfers some amount of the native cryptocurrency (bitcoin, or BTC) between parties. In Ethereum [8], transactions were extended from transferring the native cryptocurrency (ETH) to the creation or execution of smart contracts, i.e., on-chain software programs. Smart contracts can store data, hold tokens, and transfer tokens according to the contract’s internal logic, e.g., to implement an auction. A major use case for smart contracts is to define new cryptocurrencies that are different from the blockchain’s native cryptocurrency. In Ethereum, this is particularly facilitated by the ERC20 token standard [51] which defines a common set of smart contract functions required by any new cryptocurrency. Smart contracts can also support NFTs, e.g., via OpenSea or Mintable.
@@ 2.1 General Background
Stablecoins. A stablecoin is a cryptocurrency whose value is pegged to an underlying fiat currency, e.g., the US dollar. [^1] Cryptocurrencies whose value is pegged to a real asset such as gold, e.g., the PAXG token, are typically not considered stablecoins, e.g., they are not labelled as such by CoinMarketCap. There are multiple types of stablecoins [44, 32], of which the most common ones are:
  1. 1.
@@
    Crypto-collateral-backed stablecoins (or crypto-backed stablecoins for brevity), which are backed by other cryptocurrencies that are stored in one or more smart contracts. In this case, the guarantee is that in case of system collapse, each stablecoin can be redeemed for a basket of tokens whose value equals one unit of the pegged currency.
  3. 3.
@@
    Algorithmic stablecoins, which do not use reserves or collateral and which attempt to maintain a peg by adjusting the supply of a paired token [44].
@@
A multitude of fiat-backed stablecoins exist, the most prominent (as of June 2023) being Tether (USDT), USD Coin (USDC), and Binance USD (BUSD). The largest crypto-backed stablecoin is the Dai stablecoin (DAI), which is maintained by the MakerDAO foundation and which we discuss in more detail in Section 2.2. During our observation period (as discussed in Section 5), most fiat- and crypto-backed stablecoins did an overall sound job in maintaining their peg with the US dollar: the observed price ranges of DAI, USDT, USDC, and BUSD were [0.9921, 1.0062], [0.9834, 1.0005], [0.9991, 1.0073] and [0.9988, 1.0126] respectively. The analysis in [37] suggests that fiat-backed stablecoins have a trade-off between price stability and risk of system failure in the sense that better arbitrage improves the former at the cost of increasing the latter.
@@
Fiat-backed stablecoins introduce a trust assumption, namely on the agents who handle and store the fiat currency. In Tether, this trust assumption has been widely questioned – accusations include Tether using fiat collateral to buy high-risk assets [20] and creating unbacked stablecoins to prop up the price of bitcoins [25]. Furthermore, algorithmic coins have nothing to underpin them in case of a death spiral [44, 6], as witnessed by the loss of over 60 billion USD when the Terra/Luna stablecoin system failed. As such, in the rest of this paper we will focus on crypto-backed stablecoins such as Dai.
2.2 The Dai Stablecoin System (DSS)

### v4

@@
Abstract.
Decentralized Finance (DeFi), in which digital assets are exchanged without trusted intermediaries, has grown rapidly in value in recent years. Stablecoins, which are pegged to a non-volatile asset such as the US dollar, are a prominent feature of DeFi as they mitigate the risk associated with price fluctuations. However, existing stablecoin systems are tied to individual blockchain platforms, and trusted parties or complex protocols are needed to exchange stablecoin tokens between blockchains. Our goal is to design a practical stablecoin system for cross-chain commerce, and to do so we must overcome two main challenges. The first is to support a large and growing number of blockchains efficiently. The second is for the stablecoin to be resilient to blockchain platform failures and to price fluctuations that affect its collateral. We present CroCoDai to address these challenges. We demonstrate CroCoDai’s efficiency by comparing the performance of a prototype implementation to related baselines, and its resilience through an empirical analysis of historical token price data.
1. Introduction
Decentralized Finance (DeFi) promises to revolutionize finance by removing the need for trusted middlemen to process transactions. In DeFi, transactions commonly involve a change of ownership of digital tokens that exist on a distributed ledger such as a blockchain [^1] In the remainder of this paper, we will use the terms ‘blockchain’ or ‘chain’ to refer to any type of distributed ledger, including ledgers that are technically not chains (e.g., Avalanche’s directed acyclic graphs (rocket2018snowflake, )). – e.g., one user sending bitcoins to another (btc_origin, ). Many blockchain platforms (e.g., Ethereum (buterin2014next, )) also support smart contracts, which enable automated versions of financial concepts such as auctions, loans, swaps, decentralized exchanges, and ownership shares. The total value of the tokens held in DeFi smart contracts peaked at over \$150 billion (USD) in late 2021 (werner2021sok, ; defillama, ), with at least \$100 billion held in decentralized exchanges alone (xu2023sok, ), however it had fallen to around \$50 billion as of December 2023 (defillama, ).
(a) Fully Siloed
(b) Wrapped + DAI
(c) CroCoDai
Figure 1. Overview of different cross-chain stablecoin systems. V, C, and G respectively indicate the presence of vault, stablecoin, and governance contracts.
The DeFi ecosystem is currently fragmented over a multitude of different blockchain platforms which each offer their own choices in terms of efficiency, security, functionality, and structure of DeFi contracts. For example, Cardano (hoskinson2017cardano, ; kiayias2017ouroboros, ), Solana (yakovenko2018solana, ), Avalanche (rocket2019scalable, ), and Algorand (gilad2017algorand, ) each had a valuation of more than a billion USD in early 2023. These blockchains are not designed to interoperate: i.e., there are no native mechanisms for transferring tokens between chains. However, the emergence of blockchain bridges (bridges, ; celer, ) and dedicated interoperability platforms such as Cosmos (Polkadot20, ) and Polkadot (cosmos20, ) demonstrates an increasing demand for cross-chain interoperability. A major challenge in cross-chain trades is atomicity: the transactions in a trade must be committed on all blockchains, or not at all. Existing protocols for cross-chain atomicity, including relays and hashed time locks (herlihy2021cross, ; belchior2021survey, ; thyagarajan2022universal, ), essentially lock the assets involved in a trade and only release (or return) them once the transactions have been committed (or aborted) on all blockchains. While ensuring atomicity, this approach increases the financial risks for the users in the following ways: (1) the tokens held in escrow may lose value due to price fluctuations, (2) one party may abort the transaction, leading to an opportunity cost for the other parties, and (3) the DeFi platforms that facilitate the trade may be faulty or malicious . These risks are interrelated: if there is a significant change in the value of the tokens in escrow (1), then this may cause a party to abandon a trade (2), creating an opportunity cost to the other users. Price volatility can be mitigated in principle if the tokens held in escrow are stablecoins (klages2020stablecoins, ; cao2021designing, ; moin2020sok, ) whose value is pegged to a less volatile real-world asset (e.g., the US dollar). However, a flawed stablecoin design can itself put users’ tokens at risk, as witnessed by the collapse of the Terra/Luna stablecoin in May 2022 (briola2023anatomy, ), in which stablecoins worth \$currency-dollar\\$\$60 billion lost nearly all of their value.
Our goal is to design a practical stablecoin as a central point of exchange for cross-chain commerce. We focus on stablecoin designs that are backed with crypto collateral – e.g., Dai (daiwhitepaper, ) on the Ethereum blockchain – to minimize the dependence on trusted parties outside the DeFi ecosystem (e.g., banks and asset management companies). To realize our goal, we must address two main challenges. The first challenge is to support a large and growing number of blockchains efficiently, including new and relatively low-use platforms. These low-use blockchains typically have lower transaction fees, but also less liquidity and existing DeFi support. The second challenge is to ensure that the stablecoin is resilient to black swan events (e.g., 51% attacks or critical smart contract bugs) that affect individual blockchain platforms, and to price fluctuations that affect multiple blockchains simultaneously. If the total value of the collateral drops below the intended value of the issued stablecoins – which we will refer to as system failure – then this may trigger the DeFi equivalent of a bank run, forcing stablecoin holders to sell their tokens at a loss. Our key insight is that a wide range of collateral can address both challenges. In particular, it allows users from smaller chains to create robust stablecoins without interacting with a more expensive blockchain or relying on a trusted third party. Furthermore, it reduces the exposure of the stablecoin system as a whole to sudden crashes and price fluctuations.
In this work, we present CroCoDai, a stablecoin for cross-chain commerce that addresses the above challenges. Compared to its closest baseline approaches, CroCoDai has greater efficiency, i.e., lower transaction costs, and better or similar resilience to black swan events and price fluctuations.
A first baseline is for each blockchain to support its own crypto-backed stablecoin and required components: vaults to store collateral (V), coin contracts to process stablecoin transfers (C), and governance contracts for system-level decisions (G) – this is depicted in Figure 1a. In this baseline, stablecoins can be transferred between chains through bridges that are unique to each chain pair. Such bridges either lock tokens as collateral on one chain and create wrapped (wbtc, ) tokens on the other, or burn tokens (karantias2020proof, ) on one chain and mint them on the other. [^2] The Circle-created USDC bridge between Ethereum and Avalanche (usdcbridge, ) is an example of the latter. However, the lack of liquidity on smaller blockchains makes it difficult to find collateral locally, and each individual stablecoin is highly vulnerable to black swan events and price fluctuations on its platform.
A second baseline is to combine an existing crypto-backed stablecoin on a large (i.e., high-use) blockchain, e.g., Ethereum’s Dai, with two bridges for each smaller chain: one to deposit wrapped versions of the smaller chains’ native tokens on the large chain for collateral, and another to create wrapped versions of the stablecoin on the smaller chains. In this setting, each smaller chain has a vault contract (V) for collateral and an independent coin contract (C) for the wrapped stablecoins – see Figure 1b. However, this approach requires unnecessary transactions on the large blockchain to create stablecoins or move them between chains – such transactions are typically more expensive than on the smaller chains. By contrast, CroCoDai uses integrated vault and coin contracts (V and C) on each smaller chain, and leverages the governance layer on a dedicated relay chain for cross-chain transactions and system-level decisions, as depicted in Figure 1c.
In this paper, we make the following contributions:
  1. (1)
    We present CroCoDai, a stablecoin for cross-chain commerce (Section 3), and explore two main design variants: one uses a quorum of trusted nodes for the relay chain, and the other uses an existing blockchain (Section 3.3). While CroCoDai re-uses parts of the existing Dai stablecoin system, interactions with vaults are strictly local to the smaller chains, and it supports cross-chain price oracles to improve price discovery on smaller blockchains.
  2. (2)
@@ 1. Introduction
  3. (3)
[…]
In summary, CroCoDai is cheaper than a baseline with wrapped tokens, is less vulnerable to price fluctuations and black-swan events than a siloed stablecoin, and is not reliant on any trusted third parties unlike existing stablecoins such as Tether and Paypal USD. By facilitating stablecoin use on smaller chains and stablecoin transfers between chains, CroCoDai is a practical medium of exchange in cross-chain commerce (datta2024blockchain, ). To aid reproducibility, we have made the data and the source code for our experiments publicly available via https://github.com/ntublockchain/crocodai.
@@
The remainder of the paper is structured as follows. Section 2 provides background information on DeFi and related work. Section 3 presents the design of CroCoDai. We analyze CroCoDai’s resilience against black swan events in Section 4 and price fluctuations in Section 5. Section 6 details the prototype implementation and corresponding experimental results. Section 7 discusses future work and concludes the paper.
2. Background
2.1. General Background
Tokens and DeFi. In DeFi, the principal function of blockchain transactions is to transfer tokens, which represent ownership of assets. There are two common types of tokens: fungible tokens – also commonly called cryptocurrencies – and non-fungible tokens (NFTs) (das2022understanding, ). Fungible tokens of the same type are identical, whereas each NFT represents a unique asset. Smart contracts can be used to define cryptocurrencies that are different from the blockchain’s native cryptocurrency. In Ethereum, this is particularly facilitated by the ERC20 token standard (erc20, ) which defines a common set of smart contract functions required by any new cryptocurrency. Smart contracts can also support NFTs, e.g., via OpenSea or Mintable. A prominent use case for user-defined tokens is to create a version of one blockchain’s token on another blockchain – e.g., Wrapped Bitcoin (WBTC) on the Ethereum blockchain. In general, wrapped tokens (wbtc, ) are tokens maintained by custodians – e.g., for WBTC this includes the Kyber Network and BitGo – who are trusted to create wrapped tokens if and only if an equivalent amount of tokens under their control have been locked on the supporting blockchain. Xclaim (zamyatin2019xclaim, ) is an academic proposal that is similar to WBTC but which requires custodians to deposit additional collateral, thus adding a financial penalty as a disincentive for abuse (e.g., stealing the locked tokens).
@@ 2.1. General Background
Stablecoins. A stablecoin is a cryptocurrency whose value is pegged to an underlying fiat currency, e.g., the US dollar. There are multiple types of stablecoins (moin2020sok, ; klages2020stablecoins, ), of which the most common ones are:
  1. (1)
@@
    Crypto-collateral-backed stablecoins (or crypto-backed stablecoins for brevity), which are backed by other cryptocurrencies that are stored in one or more smart contracts. In this case, the guarantee is that in case of system shutdown, each stablecoin can be redeemed for a basket of tokens whose value equals one unit of the pegged currency.
  3. (3)
@@
    Algorithmic stablecoins, which do not use reserves or collateral and which attempt to maintain a peg by adjusting the supply of a paired token (moin2020sok, ).
A multitude of fiat-backed stablecoins exist, the most prominent (as of September 2023) being Tether (USDT), USD Coin (USDC), Binance USD (BUSD), and TrueUSD (TUSD). The most prominent crypto-backed stablecoin is DAI on the Ethereum blockchain, which is maintained by the MakerDAO foundation and which we discuss in more detail in Section 2.2. Other examples of crypto-backed stablecoins include Liquity USD (LUSD) and Reflexer’s RAI, which are backed entirely by ETH. Other recent stablecoin proposals include FRAX (kazemian2022frax, ), which is an algorithmic stablecoin with a cryto-backed component, and Celo (clabs2019celo, ), which is backed by a mixture of fiat and crypto assets.
@@
During our observation period (see Section 5), most fiat- and crypto-backed stablecoins did an overall sound job in maintaining their peg with the US dollar: the observed price ranges of DAI, USDT, USDC, BUSD, and TUSD were [0.9921, 1.0062], [0.9834, 1.0005], [0.9991, 1.0073], [0.9988, 1.0126], and [0.9937, 1.0056], respectively. The analysis in (ma2023stablecoin, ) suggests that fiat-backed stablecoins have a trade-off between price stability and risk of system failure in the sense that better arbitrage improves the former at the cost of increasing the latter. Recently, established financial services have launched their own fiat-backed stablecoins, e.g., PayPal’s PYUSD, which is available on the Ethereum and Solana blockchains (paypal, ). Finally, there have been proposals by central banks to issue their own digital tokens, which would have the same functionality as stablecoins – e.g., the digital euro, which recently completed its 2-year investigation phase and is now moving to the preparation phase (euro, ).
@@
Fiat-backed stablecoins introduce a trust assumption, namely on the agents who handle and store the fiat currency. In Tether, this trust assumption has been widely questioned – accusations include Tether using fiat collateral to buy high-risk assets (tetherallegation, ) and creating unbacked stablecoins to prop up the price of bitcoins (griffin2020bitcoin, ). Furthermore, algorithmic coins have nothing to underpin them in case of a death spiral (moin2020sok, ; briola2023anatomy, ), as witnessed by the loss of over 60 billion USD when the Terra/Luna stablecoin system failed. The main threat to crypto-backed stablecoins is that the system may fail if the value of the collateral decreases too rapidly – however, we show in this paper that this risk can be made small with sufficient overcollateralization. In the following, we focus particularly on Dai, which is the most prominent crypto-backed stablecoin.
2.2. The Dai Stablecoin System (DSS)